Gas systems weekly project meeting

→ Europe/Zurich
32/1-A24 (CERN)

32/1-A24

CERN

40
Show room on map
gas systems project meeting
Zoom Meeting ID
62095888225
Host
Beatrice Mandelli
Useful links
Join via phone
Zoom URL

Gas systems 2026-08-27 09:00(GMT+2:00)

Key Outcomes

Multiple concurrent gas system issues were reviewed across ALICE, ATLAS, CMS, and LHCb detectors. The RPC circulation pumps (Pump 2 failed, Pump 3 active) and RPC Rack 70 valve problems are the two most critical open items.

For CLOUD Zone access remains restricted until at least the 21st, limiting on-site intervention. A target run start of September was noted but may be postponed.

 

Incidents Resolved

·         CSC input valve, Rack 61: Chambers were disconnected without notification; gas flowed through bubbler supply, pump pressure dropped, regulation valve closed at −3.5, causing rack stop — root cause identified as miscommunication, not equipment failure.

·         RICH1 distillation cooling circuit: Misassignment of temperature sign caused failure; corrected and resolved.

·         904 Mass Flow Controller: Confirmed complete.

·         CSC panel: Work finished; glass rotameter replacement pending; ticket closed with final cost logged; TID to be completed.

·         RPC circulation pump issue: Running on Pump 1 exceeded 3,000 hours; Pump 2 failed (pressure dropped); Pump 3 activated but showed uneven membrane pressure. Opening and resealing the membrane depressurized it and restored operation — currently stable on Pump 3 for 2–3 days.

 

Active Issues

RPC Purifier 2

·         Pump not reaching vacuum; purifier stopped last week.

·         Disassembly and repair planned for current/next day; experiment confirmed oxygen increase is acceptable short-term.

RPC Circulation pump 2

·         Membrane pressure concern: All membranes showing slowly rising inter-membrane pressure since July despite no maintenance changes since January; non-return valve (metallic, not 100% tight) suspected as backflow source.

·         non-return valve last changed in 2024 (two years old); replacement under consideration.

Rack 70 RPC — CMS

·         Problem 1: Output ECONEX valve causing pressure perturbations; temporarily forced to 30% — motor/electronics suspected, replacement likely needed.

·         Problem 2: Input pressure increased after chamber repair by Alexander; unclear whether cause is the repair or a dirty manifold/supply.

·         Manifold cleaning proposed to isolate root cause; if pressure normalizes, same issue may exist in other racks.

·         Ticket opened with Maria Christina; intervention tentatively planned for mid-September.

·         Two problems treated separately: valve electronics issue is independent of the pressure increase.

 

Status Updates by Experiment

·         ALICE TPC: Switched from Neon to CO2 injection (change made without notification during holiday); Neon still present in volume.

·         ATLAS RPC purifier: Stopped; restart targeted for Monday at earliest.

·         Analysis chain: Action to review which detectors are active and whether monitoring table remains valid.

·         NA62 CEDAR detector: Maintained at 100 mbar nitrogen pressure; straw system running.

·         LHCb UT pressure regulation: Handle of Tescom regulation faulty; jira ticket submitted; Team to be informed next week. 15Need to check logic: When system stops manually, flow stops — unclear if this is intended behavior; on automatic power cut, system should switch to dry air. Action: Add exercise/drill to action list to test switchover scenarios (N2 → dry air, manual vs. automatic stop)

 

DSS Alarm Phone Call Issue

·         Control room (2231) sends generic DSS alarm calls without specifying which experiment — recipient cannot determine relevance.

·         Action: Identify and contact the person responsible for the alarm configuration to add experiment identification.

 

Access & Planning

·         CLOUD: Zone access closed until 21st; scaffolding removal and large instrument installation scheduled for two consecutive Mondays.

·         Mixer for Florian (LHCb/RICH): In planning;

·         second mixer (recirculation electrical) awaiting Jonathan's return (back this week).

·         Workshop week planning: FSU Team not fully present; ATLAS work (humidifier dismounting deferred, other Atlas work ongoing) is the priority; RPC In62 deferred to next week (half-day effort).

 

Action Items

·         Davide: Complete RPC purifier pump disassembly and repair today/tomorrow.

·         Andrea: Send CSC panel ticket link via WhatsApp; complete TID with final cost.

·         Andrea + Maria Christina: Plan Rack 70 intervention (valve replacement + manifold cleaning) for mid-September.

·         Patrick: Inform LHCb about UT pressure regulation handle issue next week.

·         Patrick: Add LHCb-UT gas system stop/switchover behaviour check to action list.

·         Louis-Philippe: Identify DSS alarm contact at control room and request experiment-specific identification.

·         Team: Review LHC analysis chain detector activity table for relevance.

·         Davide: Monitor ATLAS-RPC Pump 3 membrane pressure; evaluate gasket replacement.

 

 

 

 

 

 

 

 

 

piquet:

 

LHCb pump filters to be changed or something (for LS3)-> TICKET

 

O2+H2O analyser to be checked ->  TICKET

 

ATLAS-RPC rack 66 and 69 pressure alarm during purif 2 change over -> TICKET

 

NA62

glitch of CEDAR 6105, is it the keller? -> TICKET to check

 

ATLAS-TGC putifier Col B leak. loosing 1 bar in 1 day -> TICKET Davide check with Argon or CO2

 

CMS-DT pressure sensor PT1303 went to 0 for 20 minutes -> TICKET to follow Andrea

 

ATLAS-RPC purif -> affecting chamber pressure -> R66 check

12/2 ATL-RPC purif 1 some software issue and sms not indicating the right reason for the alarm -> TICKET

 

IR status, spares, .... --> TICKET

test binary gas analyser... -> TICKET

outcome of binary gas analyser test --> TICKET & buy a 2nd one --> Maria Cristina

DP pressure for pump filter in ALICE and LHCb -> TICKET -> @LPh + @Patrick [GASSYSTEMS-903] DP pressure for pump filter increasing - CERN Central Jira  -> Stefania

chiller fluids to uniform between CMS, LHCb and ATLAS -> TICKET Maria Cristina 

 

check if NV in purif is really useful --> TICKET

 

Calibration bottles O2-H2O status

 

ALICE

70 report from Anité [GASSYSTEMS-70] TPC gas quality - filters to Benoit - CERN Central Jira (I don't see the report, can you please add it?)

several alarms blocked: [GASSYSTEMS-628] ALICE alarms blocked - CERN Central Jira

TOF gas mixture analysis -> [GASSYSTEMS-639] TOF gas mixture analysis - CERN Central Jira & ELOG

TRD membrane is connected ? -> [GASSYSTEMS-717] TRD membrane status - CERN Central Jira

P&ID drawing check who is doing -> [GASSYSTEMS-736] ALICE P&ID drawing check - CERN Central Jira

MID swap of pumps during MD because sometime the system stop -> TICKET

 

 

 

ATLAS

MDT frequency converter (need to be replaced TICKET) ELOG - ATLAS MDT Test convertisseur 4101 avec la pompe 4201 

RPC distribution rack regulation valve (TICKET)

RPC secondary mixer (TICKET), info to be given to ATLAS

RPC purif 3 circuit breakers & meeting -> [GASSYSTEMS-718] RPC purif 3 electrical issue - CERN Central Jira --> meeting when?

ATLAS-TGC flow difference --> [GASSYSTEMS-612] ATLAS-TGC flow difference - CERN Central Jira

 

 

CMS

RPC purif test at workshop, PLC needed -> TICKET

R134a insulation - isol'tut to check and see what to do -> ticket IO [GASSYSTEMS-578] R134a insulation - isol'tut to check and see what to do - CERN Central Jira

How much gas flow we need for the RPC purifier? -->  [GASSYSTEMS-579] CMS-RPC How much gas flow we need for the RPC purifiers? - CERN Central Jira @MariaCristina notes to be added

RPC bypass pump too much closed --> [GASSYSTEMS-604] RPC bypass pump too much closed - CERN Central Jira

RPC with CO2 like ATLAS and effect on R134a recuperation -> [GASSYSTEMS-606] RPC with CO2 like ATLAS and effect on R134a recuperation - CERN Central Jira

R134a Upgrade (hardware and software) to be tested -> [GASSYSTEMS-594] R134a Upgrade (hardware and software) to be tested - CERN Central Jira

DT vaisala reading difference between SG and US -> [GASSYSTEMS-677] CMS DT vaisala reading difference between SG and US - CERN Central Jira

CMS-FGR -> TICKET should we add a module. Do we need to buy something? follow-up

CMS-FGR -> TICKET Maria Cristina già fatto per bypass

CMS-FGR -> TICKET flowmeter for Huber

CMS-RPC IR not stable -> elog Maria Cristina -> TICKET

CMS-ID flushing. refurbishing of old racks material and manpower need to be estimated -> TICKET @Gianluca ticket number for project

R134a huber temperature can be controlled by PLC -> TICKET

 

 

LHCb

TS1-> RICH1 leak search (postponed), SciFi some work from others expected

RICH1 what we learn from 1st fill -> [GASSYSTEMS-608] RICH1 what we learn from 1st fill - CERN Central Jira

summary of the C4F10 bottles --> TICKET or ELOG -> [GASSYSTEMS-660] summary of the C4F10 bottles - CERN Central Jira

alarms blocked: [GASSYSTEMS-631] LHCb alarms blocked - CERN Central Jira

RICH2 filter blocking -> [GASSYSTEMS-731] RICH2 pump filter differential pressure sensor not working - CERN Central Jira

P&ID drawing check who is doing -> [GASSYSTEMS-739] LHCb P&ID drawing check - CERN Central Jira

RICH1 cleaning , when? not before mid-september? -> TICKET 

RICH1 remove membrane rack before end of HL-LHC operation and/or change piping-> TICKET

RICH2 injection through recuperated CF4 -> TICKET

fire detection alarms reported by Pieter, more info? but not on us. Cleaning of SG room -> TICKET 

CODEX-b interlock for flammable gas detection -> Patrick

RICH1 leak repair -> December 9th tbc -> TICKET

 

NA62

CEDAR maintenance needed -> [GASSYSTEMS-609] CEDAR maintenance needed - CERN Central Jira

change of pressure sensor used for leak warning? -> [GASSYSTEMS-741] NA62-CEDAR move to keller pressure sensor? - CERN Central Jira

 

GIF++ TGC gas system and return flow: [GASSYSTEMS-624] GIF++ TGC gas system return flow - CERN Central Jira

[GASSYSTEMS-674] MFCs to be checked - CERN Central Jira :

 

non-LHC safety valve --> check lista di Fabio

[GASSYSTEMS-172] non-LHC safety valves to be verified - CERN Central Jira

ELOG - safety valves to be verified for non-LHC systems

 

 

 

Additional requests/projects/discussions

887 CEDAR new "mixers" --> [GASSYSTEMS-509] CEDAR for beam lines - CERN Central Jira

 

AMBER

EDMS leak test? 

test certificate x-ray?

ATEX certificate?

who is technical coordinator

resource manager Oleg D?

no money? should we stop? and resume in 2026? 

Commissioning when? -> TICKET 

when they would like to start? Moritz? -> ask

 

LEAK DETECTORS LIST -> TICKET

MAINTENANCE INTERVENTIONS WORK ORDER FOR STATISTICS -> TICKET

 

PROJECT 2025

summary TICKET

 

 

examples of topics for the common discussions:

 

-cms id flushing -> Andrea

-cms r134a operation and spare -> Andrea

-cms cf4 operation and spare -> Andrea

-atl tgc nc5h12 -> Kacper

-Analysis boxes -> LPh

-C4F10 new recuperation installation -> Patrick

 

There are minutes attached to this event. Show them.
    • 09:15 → 09:25
      Piquet interventions 10m
      Speakers: Gianluca Rigoletti (CERN), persons on piquet
    • 09:25 → 09:35
      gas systems status 10m
    • 09:35 → 09:55
      news for the team
      • 09:35
        ALICE news for the team 5m
        Speakers: Louis-Philippe De Menezes (CERN), Miguel Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
      • 09:40
        ATLAS news for the team 5m
        Speakers: Davide Casarotto (CERN), Maximilian Lukas Kerker (Rheinisch Westfaelische Tech. Hoch. (DE))

        RPC mixture changed: going to 90% CO2 , 5% isob , 5% R134A

      • 09:45
        CMS news for the team 5m
        Speakers: Andrea D'Auria (CERN), Maria Cristina Arena (CERN)

        LS3 officially started for CMS.

        Cavern closed until 29 of June but all the gas systems already in LS configuration (no purif, no flammable and F gases)

      • 09:50
        LHCb and CODEX-b news for the team 5m
        Speakers: Patrick Carrie (CERN), Stefania Juks (Université Paris-Saclay (FR))

        CODEX purge 24/6

        dismountling 25/6

        put inn our planning dismountling gas recirculation system CODEX

    • 09:55 → 10:05
      Review new ticket or updated during current week 10m

      Jira dashboard: https://its.cern.ch/jira/secure/RapidBoard.jspa?rapidView=9098&view=detail
      Confluence page: https://confluence.cern.ch/display/KBFGS

      Speaker: Roberto Guida (CERN)
      • Try to order a new filter
      • Consider replacing filters with filters from CMS everywhere
    • 10:05 → 10:10
      other M&O news 5m
      Speakers: Beatrice Mandelli (CERN), Gianluca Rigoletti (CERN)
    • 10:10 → 10:15
      Gas mixture analysis for LHC experiments 5m

      https://epdt-gasmonitoring.web.cern.ch/

      Speaker: Maria Cristina Arena (CERN)
    • 10:15 → 10:25
      projects ready for workshop (and to be planned) 10m
      Speakers: Albin Mekkattukulam Joshy (University of Applied Sciences Bremerhaven (AF)), Amin Bouzaiene (EPFL - Ecole Polytechnique Federale Lausanne (CH)), Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Davide Casarotto (CERN), Fabio Bordon, Gianluca Rigoletti (CERN), Louis-Philippe De Menezes (CERN), Maria Cristina Arena (CERN), Maximilian Lukas Kerker (Rheinisch Westfaelische Tech. Hoch. (DE)), Miguel Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT)), Patrick Carrie (CERN), Roberto Guida (CERN), Stefania Juks (Université Paris-Saclay (FR))

       

      -mixer for Florian (Gianluca)

      -2nd micro recirculation - electrical (Stefania)

      -next week Fabio at GIF Monday Sept 7h

    • 10:25 → 10:35
      CLOUD 10m
      Speaker: Louis-Philippe De Menezes (CERN)
    • 10:35 → 10:40
      FSU planning for coming weeks & Jobs status 5m
      Speaker: Andrea D'Auria (CERN)
    • 10:40 → 11:00
      EAM/Infor: update
      Convener: Miguel Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
      • 10:40
        interventions from past week 5m
        Speakers: Emiliano Vaccaro (Politecnico di Torino (IT)), Miguel Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))

        -interventions per type (preventive, corrective, ...)

        -interventions per module

        -intervention per component

        -part used during intervention

        -time spent on intervention and total

        -divide between working hours and piquet

         

        -develop problem code list

      • 10:45
        maintenance and spare parts follow-up 5m
        Speaker: Miguel Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
      • 10:50
        pump working hours 5m
        Speaker: Miguel Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT))
    • 11:00 → 11:05
      LS3 planning preparation 5m
    • 11:05 → 11:15
      aob and requests for jobs at workshop 10m
      Speakers: Albin Mekkattukulam Joshy (University of Applied Sciences Bremerhaven (AF)), Amin Bouzaiene (EPFL - Ecole Polytechnique Federale Lausanne (CH)), Andrea D'Auria (CERN), Andrea Silvan (Universita degli Studi di Trento and INFN (IT)), Beatrice Mandelli (CERN), Bjorn Pascal Wagner, Clea Fourreaux (Université de Savoie Mont Blanc (FR)), Davide Casarotto (CERN), Emiliano Vaccaro (Politecnico di Torino (IT)), Fabio Bordon, Gianluca Rigoletti (CERN), Henrietta Szalai (Ludwig Maximilians Universitat (DE)), Louis-Philippe De Menezes (CERN), Maria Cristina Arena (CERN), Maximilian Lukas Kerker (Rheinisch Westfaelische Tech. Hoch. (DE)), Michael Vogl (Fachhochschule Regensburg (DE)), Miguel Afonso (Laboratory of Instrumentation and Experimental Particle Physics (PT)), Olaoluwa Emmanuel Olaofe (Aston University (GB)), Patrick Carrie (CERN), Roberto Guida (CERN), Ryan Oatley (University of Leeds (GB)), Stefania Juks (Université Paris-Saclay (FR)), Thomas Kattikanayil Thomas (Technical University of Liberec (CZ))
    • 11:15 → 11:20
      Henrietta Szalai 5m
      Speaker: Henrietta Szalai (Ludwig Maximilians Universitat (DE))
    • 11:20 → 11:25
      Michael Vogl 5m
      Speaker: Michael Vogl (Fachhochschule Regensburg (DE))